OpenStudy (anonymous):

what is the oblique asymptote of the funtion f(x)=(x^2-5x+6)/(x+4)

OpenStudy (anonymous):

I got y=x-1 is that right?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

dividing (x^2-5x+6)/(x+4) by (x+4) yields x - 9 + 36/(x + 4), no? So your oblique asymptote is y = x - 9, no?

OpenStudy (anonymous):

Your right thank you. =D

OpenStudy (anonymous):

can you help me with this one too, what are the zeros of the function f(x)=3x^2-9x+6/3x-3

OpenStudy (anonymous):

I got 1

OpenStudy (anonymous):

You do realize that a zero of the polynomial is the same thing as an x-intercept. You are being asked to find all values of x where the function crosses the x-axis, that is where the value of the function is zero. So you want to solve for x, when f(x) = 0. f(x)=3x^2-9x+6/3x-3 = 0 In other words, solve the equation 3x^2 - 9x + 6/(3x - 3) = 0. Can you do that?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

or is it 2 or -2........IDK

OpenStudy (mertsj):

The zeros of the function occur when y is 0. So you can write: \[\frac{3x^2-9x+6}{3x-3}=0\]

OpenStudy (mertsj):

However you then realize that a fraction is 0 ONLY IF its numerator is 0 so you write: \[3x^2-9x+6=0\]

OpenStudy (mertsj):

Solve that by factoring out the 3: \[3(x^2-3x+2)=0\]

OpenStudy (mertsj):

And then factor the trinomial: \[3(x-2)(x-1)=0\]

OpenStudy (mertsj):

Set each factor equal to 0 and solve. Obviously 3 cannot be 0 so x=2 and x = 1 are zeros of the function.

OpenStudy (anonymous):

It looked like your equation was f(x)=3x^2-9x+6/(3x-3). I guess you meant f(x)=(3x^2-9x+6)/(3x-3). In that case, Mertsj has given the correct answer.